Seven and a half years ago, Chris invested $10,000 in a retirement fund that grew at the rate of 9.22%/year compounded quarterly. What is his account worth today? (Round your answer to the nearest cent.)
Answer by stanbon(75887)

Seven and a half years ago, Chris invested $10,000 in a retirement fund that grew at the rate of 9.22%/year compounded quarterly. What is his account worth today?
------------------------------
A(t) = P(1+(r/n))^(nt)
------
A(7.5) = 10,000(1+(0.0922/4))^(4*7.5)
---------
A(7.5) = 10000(1.02305)^(30)
---
A(7.5) = 10000(1.981097) = $19,810.97
========================================
Cheers,
Stan H.